OpenStudy (anonymous):

how do you simplify (3-5i)(3+5i)+(4-7i)/i+1?

OpenStudy (anonymous):

Simplify means break it down by comparing like terms. So compare like terms. Its easier than it looks. It just seems hard because you have a massive problem but ill help you

OpenStudy (anonymous):

Ok so first you want to make a problem with your like terms. I made -5i + 5i - 7i / i. Try and solve that first.

OpenStudy (anonymous):

You want to use order of operations. So going in order, -7i divided by i will equal a -7i ofcourse so now you have -5i + 5i - 7i. Negative and positive 5i will equal to 0. So you`re left with -7i. So you have one part of the problem done.

OpenStudy (anonymous):

The other problem will be 3 • 3 + 4 +1 = 14

OpenStudy (anonymous):

So my final answer is 14 - 7i

OpenStudy (anonymous):

Hope I helped :)

OpenStudy (anonymous):

We know that the formula (a - b)*(a+b) = a^2 - b^2 so use this for ( 3 - 5i) * (3+5i) = 3^2 - 5i^2 = 9 - 25i^2 since i^2 = -1 we will get 9 - (-25) = 9 + 25 = 34 so our problem become 34 +4 - 7i / i+1 = ( 38 - 7i )/(i+1)

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

You are welcome :)